OpenStudy (anonymous):

Type the second number for each pair such that square root is 6 is the geometric mean for the pair. What is the second number in the pair if the first number is 2?

OpenStudy (anonymous):

The geometric mean of n positive numbers is the nth root of their product. For example, the geometric mean of 3 and 12 is √(3*12) = √36 = 6. I think what you’re asking is: 2 * x = √6 x = 3 because 2 * 3 = 6, and thus the geometric mean is √6. Hope that helps!